If A and B are independent events with P.A/ D 0:35 and P.B/ D 0:55, then P.AjB/ is 0:35=0:55 D

0:64:


1
Expert's answer
2021-06-21T13:44:47-0400

If "A" and "B" are independent events, then


"P(A\cap B)=P(A)P(B)"

Hence

"P(A|B)=\dfrac{P(A\cap B)}{P(B)}=\dfrac{P(A)P(B)}{P(B)}=P(A)"

"P(A|B)=0.35"